Shift Summary for Owl, Sept.10, 1999

Angela Bellavance (bellavan@rice.edu)
Fri, 10 Sep 1999 07:11:02 -0500 (CDT)

Shift Summary for owl shift, 10/Sep/99

On Shift: S.Taegar, A.Bellavance

Hours of beam delivered at all intensities: 4.9/7.0
Hours of beam delivered at nominal intensity: 4.9
Hours of E832 data (special runs not included): 4.7
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: difficulty initing run, trigger errors

========================================================================
Runs ended this shift:
Run 14488 - 13M events - BEAM_832
Run 14489-90 - Flops
Run 14491 - ongoing - BEAM_832

Tapes brought to FCC this shift: QKB144,
QKC408-410,
QKK408-410,
QKS408-410
(actually this mission was entrusted to the day shift)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: E832 datataking, fixing the DAQ, 832 tape labelling,
thumb twiddling,...

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14488 119 00:00 Taking data. Muon run already done for tonight.
124 Beam gone one spill.
02:48 See problem with kq# and ban# kumacs. Page JasonL
who says he'll look into it. (Red 2X-3 Bn and
2X-1 Bn and frowny on spottiness plot).
Also note that lsr# and fpedcal# have no reference
plots in the book.
263 03:14 Beam goes away after getting worse over the last
~10 spills.
03:20 Ch.13: Booster injection probs. Est 15 mins.
03:46 Ch.13: Booster injection power supply failure.
Est. 1 hour.
03:49 Tapes are 83.6% full. Decide to stop run and load
fresh tapes.
14489 Run init fails because sleepy Angela forgot to
close a tape drive door. Do a drt_fresh_start.
14490 Run hangs during "start"; not sure of problem.
Do a drt_fresh_start.
04:54 Beam returns and we're not ready! Argh!
05:02 Abort at A2. Beam goes away.
14491 1 05:09 Start run, but now no beam.
05:18 Ch.13: TeV quench protection system problem.
Est. 20 minutes.
Meanwhile: Triggers hang because of 3X KQ DYC Busy ->
reloaded L2Proc:Kumquat. Then they hang with a
crate 53 AZLAT busy -> replug command cable on
UCLA scaler which had req light on. Then pipeline
errors -> reset pipeline.
05:34 We are back alive! (And just in time :)
20 05:35 Beam returns!!
22-23 Beam gone two spills.
46 Ktevana errors - KFRENDs.
56 06:24 Ktevana errors - KFRENDs.
76-77 06:50 Beam gone a couple spills. Greg shows up early
for his day shift. Wow!
07:07 Page the absent Arun; he's on his way.
07:10 Ed shows up. We decide to hand off to him and
Greg.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Whoever is in charge of "lsr" and "fpedcal" plots needs to get them into
the newest reference plotbook.

For Sydney, Angela

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke